OpenStudy (anonymous):

Determine the solutions to the quadratic equation -4x2 - 32x - 64 = 0. How do you work this kind of question.?

OpenStudy (anonymous):

simplifyin v get x^2+8+16=0 thus d factors r x+4.x+4

OpenStudy (anonymous):

I don't understand how to factor.

OpenStudy (anonymous):

after dividing the entire equation by -4 ,2 simplify it, u need 2 check out d product of d coefficient of x^2 nd the constant term.in dis case it is 16.u need to find out certain values whose product is 16 while the sum is the coefficient of x which is 8 in dis case .the values which satisfy this condition are 4 nd 4! now u need 2 replace 8x by 4x+4x.the equation becums x^2+4x+4x+16=0 solvin it v get (x+4)+4(x+4)=0 or (x+4)(x+4)

OpenStudy (anonymous):

aply quardatic formula to solve this question

OpenStudy (anonymous):

in this a=-4 b=-32 c=-64